OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. (x^3/4)^3/7

OpenStudy (anonymous):

x^9/28

OpenStudy (anonymous):

power raised to a power= multiply powers. eg. (2^5)^2=2^10 power times a power= add powers. eg. (2^5)(2^2)=2^7